Show that A = B where A = {x│x² - 4x +4 =1}, B = {1,3}


1
Expert's answer
2021-11-15T20:23:21-0500

Solution:

A = {x│x² - 4x +4 =1}

"x^2-4x+4=1
\\\Rightarrow x^2-4x+3=0
\\\Rightarrow x^2-3x-x+3=0
\\\Rightarrow (x-3)(x-1)=0
\\\Rightarrow x-3=0, x-1=0
\\\Rightarrow x=3, x=1"

Thus, A = {3,1}

And B = {1,3} is given

So, A = B